This is the problem I am stuck on.
5(y-2)=-2(12-9y)+y
This is what I did:
5(y-2)=-2(12-9y)+y
5y-10=-24+18y+y
5y-10=-24+19y
5y-10+10=(-24+10)+19y
5y=14+19y
5y-19y=14
-14y=14
y=-1
Is this right or did I mess up somewhere.
Thanks for any help.

Answer by Earlsdon(6103)

Yes, I'm afraid so! You did fine up to and including the fourth step of your solution, where you arrived at:
5y-10+10 = (-24+10)+19y But -24 + 10 = -14, not 14 which is what you show.
So, let's continue from there:
5y = -14 + 19y
5y-19y = -14
-14y = -14
y = 1
Check:
5(y-2) = -2(12-9y)+y
5(1-2) = -2(12-9(1))+1
5(-1) = -2(3)+ 1
-5 = -6+1
-5 = -5 Check is ok.